OpenStudy (preetha):

How do you calculate the volume of a sphere with radius 3inches in liters?

OpenStudy (amistre64):

is there any particular reason why the post turns blue?

OpenStudy (anonymous):

Asking for QH I think

OpenStudy (preetha):

Amistre is it blue on the left pane for you?

OpenStudy (amistre64):

yes it is

OpenStudy (nincompoop):

volume of sphere in general is \(V = \frac{4}{3} \pi r^3 \)

OpenStudy (preetha):

Ok, that is because I paid 10 Owlbucks to get it answered.

OpenStudy (amistre64):

you paid to get it answered? or responded to?

OpenStudy (preetha):

Well, I paid to get a Qualified Helper to answer it.

OpenStudy (nincompoop):

everything else is constant so all you need to do is multiply the value of the radius to itself three times and then multiply it by 4/3 then add the pi symbol

OpenStudy (amistre64):

is it spose to be blue for all to see? or was that intended for special people?

OpenStudy (preetha):

Blue for all to see.

OpenStudy (acxbox22):

as of now what do QH's do with owlbucks they receive as payments?

OpenStudy (amistre64):

making sure :) carry on ...

OpenStudy (preetha):

So I paid 10 OwlBucks. Nincompoop is the only QH online, he came to help. I liked his response. I rated him. He gets a rating and a share of the OB.

OpenStudy (preetha):

Anyone who wants to be a QH - read What's New and the blog.

OpenStudy (preetha):

What about the liters bit?

OpenStudy (rational):

\[\large \text{ L} = \dfrac{\text{in}^3}{61.024}\]

OpenStudy (nincompoop):

let us do a dimensional analysis

OpenStudy (rational):

you may use that to convert volume to liters in the end

OpenStudy (preetha):

Thanks Nincl Awesome job!

OpenStudy (preetha):

Thanks all of you. You were really helpful

OpenStudy (nincompoop):

I guess rational gave the conversion unit already

OpenStudy (michele_laino):

We have: 1 inches = 2.54 cm so: 3 inches = 7.62 cm or: 3 inches = 0.762 dm then the requested volume, in liters, is: V = (4*3.14/3) * (0.762)^3 of course I'm referring to liters of water

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @Preetha So I paid 10 OwlBucks. Nincompoop is the only QH online, he came to help. I liked his response. I rated him. He gets a rating and a share of the OB. \(\color{blue}{\text{End of Quote}}\) What does he do with the share of the OwlBuck?